A 2.07 gram sample of an unknown gas is found to occupy a volume of 1.57 L at a pressure of 786 mmHg and a temperature of 32 °C. Assume ideal behavior.

The molar mass of the unknown gas is
g/mol.
1
Expert's answer
2019-02-25T07:46:19-0500

We should use Ideal Gas Law to find the molar mass

PV = nRT

As n=m/M, then PV = (m/M)*RT

M = mRT/PV = 2.07 g *62.364 L*mmHg/mol*K * (273.15 +32) / (786 mmHg * 1.57 L) = 31.9 g/mol

Answer: 31.9 g/mol